OpenStudy (anonymous):

Explain, in complete sentences, how to factor 4x^2 – 4x – 48. Explain how your factors could be checked for accuracy.

OpenStudy (anonymous):

it is often easier to work with polynomials that are of the for x^2 + bx +c so we start our solution process by dividing the polynomial by 4 to obtain x^2 -x -12 we have a=1, b=-1 and c=-12 write down all possible ways to factor c in this case it would be -12 , 1 12 , -1 -6, 2 6 , -2 -4, 3 4, -3 look for the factors which add to be equal to the value of b in this case -1 -12 + 1 = -11 NOPE 12 + (-1) = 11 NOPE -6 +2 = -4 NOPE 6 + (-2) = 4 NOPE -4 + 3 = -1 YES the solution is (x-4)(x+3) the solution can be verified (x-4)(x+3) which would give us the original polynomial